# The Faiss Library

### v2

@@ 3.3 Tradeoffs
Most often only a subset of metrics matter. For example, when a very large number of searches are performed on a fixed index, index building time does not matter. Or when the number of vectors is so small that the raw database fits in RAM multiple times, then memory usage does not matter. We call the metrics that we care about the active constraints. Note that accuracy is always an active constraint because if it does not matter, an index that returns random results would be sufficient (Faiss does actually provide an IndexRandom used in some benchmarking tasks).
In the following sections, we will most often consider that the active constraints are speed, memory usage and accuracy. Thus we can for example fix a memory budget and measure the speed and accuracy of several index types and hyperparameter settings.
3.4 Exploring search-time settings
For a fixed index, there are often one or several search-time hyper-parameters that shift the tradeoff between speed and accuracy. This is for example the nprobe hyperparameter for an IndexIVF, see Section 5. In general, we define hyperparameters as discrete scalar values such that when the value is higher, the speed decreases and the accuracy increases. We can then keep only the Pareto-optimal settings, defined as settings that are the fastest for a given accuracy, or equivalently that have the highest accuracy for a given time budget [Sun et al., 2023a].
@@ 3.4 Exploring search-time settings
There is an interesting optimization in this case, which enables efficient pruning. We note a tuple of $n$ hyper-parameters $\pi=(p_{1},...,p_{n})\in\mathcal{P}=\mathcal{P}_{1}\times...\times\mathcal{P}_{n}$, and $\leq$ a partial ordering on $\mathcal{P}$: $(p_{1},..,p_{n})\leq(p^{\prime}_{1},..,p^{\prime}_{n})\Leftrightarrow\forall i,p_{i}\leq p^{\prime}_{i}$. Let $S(\pi)$ and $A(\pi)$ be the speed and accuracy obtained with this tuple of parameters. $\mathcal{P}^{*}\subset\mathcal{P}$ is the set of Pareto-optimal settings:
$\mathcal{P}^{*}=\big{\{}\pi\in\mathcal{P}|\nexists\pi^{\prime}\in\mathcal{P}\mathrm{\ s.t.\ }(S(\pi^{\prime}),A(\pi^{\prime}))>(S(\pi),A(\pi))\big{\}}$ (5)
Since the individual parameters have a monotonic effect on speed and accuracy, we know:
$\pi^{\prime}\geq\pi\Rightarrow\left\{\begin{array}[]{c}S(\pi^{\prime})\leq S(\pi)\\ A(\pi^{\prime})\geq A(\pi)\\ \end{array}\right.$ (6)
@@
$S(\pi)$$\leq$ $\widehat{S}(\pi)=\underset{\pi^{\prime}\in\widehat{\mathcal{P}}\mathrm{\ s.t.\ }\pi^{\prime}\leq\pi}{\mathrm{Inf}}S(\pi^{\prime})$ (7)
$A(\pi)$$\leq$ $\widehat{A}(\pi)=\underset{\pi^{\prime}\in\widehat{\mathcal{P}}\mathrm{\ s.t.\ }\pi^{\prime}\geq\pi}{\mathrm{Inf}}A(\pi^{\prime})$ (8)
@@
$\exists\pi^{\prime}\in\widehat{\mathcal{P}}\mathrm{\ s.t.\ }(S(\pi^{\prime}),A(\pi^{\prime}))>(\widehat{S}(\pi),\widehat{A}(\pi))\Rightarrow\pi\notin\mathcal{P}^{*}$ (9)
In practice, we evaluate settings from $\mathcal{P}$ in a random order. The pruning becomes more and more effective throughout the process. It is also more effective when the number of parameters is larger. Figure 1 shows an example with $|\mathcal{P}|=5808$ combined parameter settings. The pruning from (9) reduces this to 398 experiments, out of which $|\mathcal{P}^{*}|=87$ are optimal. The Faiss object OperatingPoints implements this.
@@ 3.5 Refining (IndexRefine)
It is possible to combine a fast and inaccurate indexing method with a slower and more accurate search [Jégou et al., 2011b, Subramanya et al., 2019, Guo et al., 2020]. This is done by querying the fast index to retrieve a shortlist of results. The more accurate search then computes more accurate search results only for the shortlist. This requires the accurate index allow efficient random access to database vectors. Some implementations use a slower storage (e.g. flash) for the second index [Subramanya et al., 2019, Sun et al., 2023b].
For the first-level index, the relevant accuracy metric is the recall at the rank that will be used for re-ranking. The 1-recall@1000 can thus sometimes be a relevant metric, even if the end application does not use the 1000th neighbor at all.
Several methods based on this refining principle do not use two separate indexes. Instead, they use two ways of interpreting the same compressed vectors: a fast and inaccurate decoding and a slower but more accurate decoding [Douze et al., 2016, Douze et al., 2018, Morozov and Babenko, 2019, Amara et al., 2022, Huijben et al., 2024] are based on this principle. The polysemous codes method [Douze et al., 2016] is implemented in Faiss’s IndexIVFPQ.
@@ 4 Compression levels
$n=\underset{i=1..N}{\mathrm{argmin}}\|q-D(C(x_{i}))\|=\underset{i=1..N}{\mathrm{argmin}}\|q-D(C_{i})\|$ (10)
Where the codes $C_{i}=C(x_{i})$ are precomputed and stored in the index. This is the asymmetric distance computation (ADC) [Jégou et al., 2010]. The symmetric distance computation (SDC) corresponds to the case when the query vector is also compressed:
$n=\underset{i=1..N}{\mathrm{argmin}}\|D(C(q))-D(C_{i})\|$ (11)
@@ 4.1 The vector codecs
The ideal vector quantizer minimizes the MSE between the original and the decompressed vectors. This is formalized in the Lloyd necessary conditions for the optimality of a quantizer [Lloyd, 1982].
@@
A very classical and simple scalar quantizer is LSH (IndexLSH), where each vector component is encoded in a single bit by comparing it to a threshold. The threshold can be set to, 0 or trained. Faiss supports efficient SDC search of binary vectors via the IndexBinary objects, see Section 4.5.
The Faiss ScalarQuantizer also supports uniform quantizers that encode a vector component into 8, 6 or 4 bits – referred to as SQ8, SQ6, SQ4. A scale and offset determine which values are reconstructed. They can be set separately for each dimension on the whole vector. The IndexRowwiseMinMax stores vectors with per-vector normalizing coefficients. Lower-precision 16-bit floating point representations are also considered as scalar quantizers, SQfp16 and SQbf16.
@@
The product quantizer (ProductQuantizer, also noted PQ) is a simple MCQ that splits the input vector into $M$ sub-vectors and quantizes them separately [Jégou et al., 2010] with a k-means quantizer. At reconstruction time, the individual reconstructions are concatenated to produce the final code. In the following, we will use the notation PQ6x10 for a product quantizer with 6 sub-vectors each encoded in 10 bits ($M=6$, $K=2^{10}$).
Additive quantizers are a family of MCQ where the reconstructions from sub-quantizers are summed up together. Finding the optimal encoding for a vector given the codebooks is NP-hard, so practical additive quantizers are heuristics to find near-optimal codes.
Faiss supports two types of additive quantizers. The residual quantizer (ResidualQuantizer) proceeds sequentially, by encoding the difference (residual) of the vector to encode and the one that is reconstructed by the previous sub-quantizers [Chen et al., 2010]. The local search quantizer (LocalSearchQuantizer) starts from a sub-optimal encoding of the vector and locally explores neighbording codes in a simulated annealing process [Martinez et al., 2016, Martinez et al., 2018]. We use notations LSQ6x10 and RQ6x10 to refer to additive quantizers with 6 codebooks of size $2^{10}$.
Faiss also supports a combination of PQ and additive quantizer, ProductResidualQuantizer. In that case, the vector is split in sub-vectors that are encoded independently with additive quantizers [Babenko and Lempitsky, 2015]. The codes from the sub-quantizers are concatenated. We use the notation PRQ2x6x10 to indicate that vectors are split in 2 and encoded independently with RQ6x10, yielding a total of 12 codebooks of size $2^{10}$.
Figure 2: The hierarchy of quantizers. Each quantizer can represent the set of reproduction values of the enclosed quantizers.
@@
Although this is not by design, it turns out that there is a strict ordering between the quantizers described before. This means that quantizer $i+1$ can have the same set of reproduction values as quantizer $i$: it is more flexible and more data adaptive. The hierarchy of quantizers is shown in Figure 2:
@@
    the scalar quantizer can be represented as a product quantizer with 1 dimension per sub-vector and uniform per-dimension quantizer;
@@
    the product quantizer can be represented as a product-additive quantizer where the additive quantizer has a single level;
@@
    the product additive quantizer is an additive quantizer where within each codebook all components outside one sub-vector are set to 0 [Babenko and Lempitsky, 2014];
@@
    the additive quantizer (and any other quantizer) can be represented as a vector quantizer where the codebook entries are the explicit enumeration of all possible reconstructions.
The implications of this hierarchy are (1) the degrees of freedom for the reproduction values of quantizer $i+1$ are larger than for $i$, so it is more accurate (2) quantizer $i+1$ has a higher capacity so it consumes more resources in terms of training time and storage overhead than $i$. In practice, the product quantizer often offers a good tradeoff, which explains its early adoption.
Figure 3: Comparison of additive quantizers in terms of encoding time vs. accuracy (MSE). Lower values are better for both. We consider two different regimes: Deep1M (low-dimensional) to 8-bytes codes and Contriever1M (high dimensional) to 64-byte codes. For some RQ variants, we indicate the beam size setting at which that tradeoff was obtained.
@@ 4.2 Vector preprocessing
To take advantage of some quantizers, it is beneficial to apply transformations to the input vectors prior to encoding them. Many of these transformations are $d$-dimensional rotations, that do not change comparison metrics like cosine, L2 and inner product.
Scalar quantizers assign the same number of bits per vector component. However, for distance comparisons, if specific vector components have a higher variance, they have more impact on the distances. In other works, a variable number of bits are assigned per component [Sandhawalia and Jégou, 2010]. However, it is simpler to apply a random rotation to the input vectors, which in Faiss can be done with a RandomRotationMatrix. The random rotation spreads the variance over all the dimensions without changing the measured distances.
An important transform is the Principal Component Analysis (PCA), that reduces the number of dimensions $d$ of the input vectors to a user-specified $d^{\prime}$. This operation (PCAMatrix) is the orthogonal linear mapping that best preserves the variance of the input distribution. It is often beneficial to apply a PCA to large input vectors before quantizing them as k-means quantizers are more likely to “fall” in local minima in high-dimensional spaces [Liu et al., 2015, Jégou et al., 2011a].
The OPQ transformation [Ge et al., 2013] is a rotation of the input space that decorrelates the distribution of each sub-vector of a product quantizer [^3] In Faiss terms, OPQ and ITQ are preprocessings. The actual quantization is performed by a subsequent product quantizer or binarization step.. This makes PQ more accurate in the case where the variance of the data is concentrated on a few components. The Faiss implementation OPQMatrix combines OPQ with a dimensionality reduction. The ITQ transformation [Gong et al., 2012] similarly rotates the input space prior to binarization (ITQMatrix).
@@ 4.3 Faiss additive quantization options
We look in more detail into the additive quantizer variants: the residual quantizer and local search quantizer. They are more complex than most quantizers because the index building time has to be considered. In fact, the accuracy of an additive quantizer of a certain size can always be increased at the cost of an increased encoding time (and training time).
Additive quantizers rely on $M$ codebooks $T_{1},...T_{M}$ of size $K$ in dimension $d$. The decoding of code $C(x)=(c_{1},...,c_{M})$ is
$x^{\prime}=D(C(x))=T_{1}[c_{1}]+...+T_{M}[c_{M}]$ (12)
Thus, decoding is unambiguous. However, there is no practical way to do exact encoding, let alone training the codebooks. Enumerating all possible encodings is of exponential complexity in $M$.
Deep1M Contriever1M
Refer to caption Refer to caption
Figure 4: Tradeoff of accuracy vs. code size for different codecs on the Deep1M and Contriever1M datasets. We show Pareto-optimal variants with larger dots and indicate the quantizer in text for some of them. Note that contriever vectors can be encoded to MSE=$2\cdot 10^{-4}$ in 768 bytes with SQ8 (that setting is widely out-of-range for the plot).
@@
RQ encoding is sequential. At stage $m$ of the encoding of $x$, RQ picks the entry that best reconstructs the residual of $x$ w.r.t. the previous encoding steps:
$c_{m}=\underset{j=1..K}{\mathrm{argmin}}\left\|\sum_{i=1}^{m-1}T_{i}[c_{i}]+T_{m}[j]-x\right\|^{2}$ (13)
This greedy approach tends to get trapped in local minima. As a mitigation, the encoder maintains a beam of max_beam_size of possible codes and picks the best code at stage $M$. This parameter adjusts the tradeoff between encoding time and accuracy.
@@
At encoding time, LSQ starts from a suboptimal encoding of the vector and proceeds with a simulated annealing optimization to refine the codes. In each iteration of the optimization step, LSQ randomly flips codes and then uses Iterated Conditional Mode (ICM) to optimize the new encoding. The number of optimization steps is set with encode_ils_iters. The LSQ codebooks are trained via an expectation-maximization procedure (similar to k-means).
@@
$\langle q,x^{\prime}\rangle=\sum_{m=1}^{M}\langle T_{m}[c_{m}],q\rangle=\sum_{m=1}^{M}\mathrm{LUT}_{m}[c_{m}]$ (14)
The lookup tables $\mathrm{LUT}_{m}$ are computed when a query vector comes in, similar to product quantizer search [Jégou et al., 2010].
This decomposition does not work to compute L2 distances. As a workaround, Faiss uses the decomposition [Babenko and Lempitsky, 2014]
$\|q-x^{\prime}\|^{2}=\|q\|^{2}+\|x^{\prime}\|^{2}-2\langle q,x^{\prime}\rangle$ (15)
Thus, the term $\|x^{\prime}\|^{2}$ must be available at search time. Depending on the setting of AdditiveQuantizer.search_type it can be appended in the stored code (ST_norm_float32), possibly compressed (ST_norm_qint8, ST_norm_qint4,…). It cal also be computed on-the-fly (ST_norm_from_LUT) with
$\|x^{\prime}\|^{2}=2\sum_{m=1}^{M}\sum_{\ell=1}^{m-1}\langle T_{m}[c_{m}],T_{\ell}[c_{\ell}]\rangle+\sum_{m=1}^{M}\|T_{m}[c_{m}]\|^{2}$ (16)
Where the norms and dot products are stored in the same lookup tables as the one used for beam search. Therefore, it is a tradeoff between memory overhead to store codes and search time overhead.
Figure 3 shows the tradeoff between encoding time and MSE. For a given code size, it is more accurate to use a smaller number of sub-quantizers $M$ and a higher $K$. GPU encoding for LSQ does not help systematically. The LUT-based encoding of RQ is interesing for RQ/PRQ quantization when the beam size is larger. For the 64-byte regime, we observe that LSQ is not competitive with RQ. PLSQ and PRQ progressively become more competitive for larger memory budgets. They are also faster, since they operate on smaller vectors.

### v3

@@ 3.3 Tradeoffs
Most often only a subset of metrics matter. For example, when a very large number of searches are performed on a fixed index, the index building time does not matter. Or when the number of vectors is so small that the raw database fits in RAM multiple times, then the memory usage does not matter. We refer to the metrics that we care about as the active constraints. Note that accuracy is always an active constraint because if it did not matter, returning random results would be sufficient (Faiss does actually provide an IndexRandom used in some benchmarking tasks).
In the following sections, we consider that the active constraints are speed, memory usage and accuracy. As such, we measure the speed and accuracy of several index types and hyperparameter settings under a fixed memory budget.
3.4 Exploring search-time settings
@@
For a fixed index, there are often one or several search-time hyperparameters that trade off speed with accuracy. For example, the nprobe hyperparameter for an IndexIVF, see Section 5. In general, we define hyperparameters as discrete scalar values such that when the value is higher, the speed decreases and the accuracy increases. We can then keep only the Pareto-optimal settings, defined as settings that are the fastest for a given accuracy, or equivalently, that have the highest accuracy for a given time budget [80].
@@
one can leverage the monotonous nature of the hyperparameters for efficient pruning. We note a tuple of $n$ hyper-parameters $\pi=(p_{1},...,p_{n})\in\mathcal{P}=\mathcal{P}_{1}\times...\times\mathcal{P}_{n}$, and $\leq$ a partial ordering on $\mathcal{P}$: $(p_{1},..,p_{n})\leq(p^{\prime}_{1},..,p^{\prime}_{n})\Leftrightarrow\forall i,p_{i}\leq p^{\prime}_{i}$. Let $S(\pi)$ and $A(\pi)$ be the speed and accuracy obtained with this tuple of parameters. $\mathcal{P}^{*}\subset\mathcal{P}$ is the set of Pareto-optimal settings:
$\mathcal{P}^{*}=\big{\{}\pi\in\mathcal{P}|\nexists\pi^{\prime}\in\mathcal{P}\mathrm{\ s.t.\ }(S(\pi^{\prime}),A(\pi^{\prime}))>(S(\pi),A(\pi))\big{\}}.$ (5)
Since the individual parameters have a monotonic effect on speed and accuracy, we have
$\pi^{\prime}\geq\pi\Rightarrow\left\{\begin{array}[]{c}S(\pi^{\prime})\leq S(\pi),\\ A(\pi^{\prime})\geq A(\pi).\\ \end{array}\right.$ (6)
@@
$S(\pi)$$\leq$ $\widehat{S}(\pi)=\underset{\pi^{\prime}\in\widehat{\mathcal{P}}\mathrm{\ s.t.\ }\pi^{\prime}\leq\pi}{\mathrm{Inf}}S(\pi^{\prime}),$ (7)
$A(\pi)$$\leq$ $\widehat{A}(\pi)=\underset{\pi^{\prime}\in\widehat{\mathcal{P}}\mathrm{\ s.t.\ }\pi^{\prime}\geq\pi}{\mathrm{Inf}}A(\pi^{\prime}).$ (8)
@@
$\exists\pi^{\prime}\in\widehat{\mathcal{P}}\mathrm{\ s.t.\ }(S(\pi^{\prime}),A(\pi^{\prime}))>(\widehat{S}(\pi),\widehat{A}(\pi))\Rightarrow\pi\notin\mathcal{P}^{*}.$ (9)
In practice, we evaluate settings from $\mathcal{P}$ in a random order. The pruning becomes more and more effective throughout the process. It is also more effective when the number of parameters is larger. Figure 1 shows an example with $|\mathcal{P}|=5808$ combined parameter settings. The pruning from (9) reduces this to 398 experiments, out of which $|\mathcal{P}^{*}|=87$ are optimal. The Faiss OperatingPoints object implements this.
@@ 3.5 Refining (IndexRefine)
One can combine a fast but inaccurate index with a slower, more accurate search. [46, 79, 36]. This is done by querying the fast index to retrieve a shortlist of results. The more accurate search then computes more accurate results only for the shortlist. This requires the accurate index to allow efficient random access to database vectors. Some implementations use a slower storage (e.g. flash) for the second index [79, 81].
For the first-level index, the relevant accuracy metric is the recall at a rank equal to the shortlist size. Thus, 1-recall@1000 can be a relevant metric, even if the end application does not use the 1000th neighbor.
Several methods based on this refining principle do not use two separate indexes. Instead, they use two ways of interpreting the same compressed vectors: a fast and inaccurate decoding and a slower but more accurate decoding [28, 29, 61, 1, 39] are based on this principle. The polysemous codes method [28] is implemented in Faiss’s IndexIVFPQ.
@@ 4 Compression levels
$n=\underset{i=1..N}{\mathrm{argmin}}\|q-D(C(x_{i}))\|=\underset{i=1..N}{\mathrm{argmin}}\|q-D(C_{i})\|,$ (10)
where the codes $C_{i}=C(x_{i})$ are precomputed and stored in the index. This is the asymmetric distance computation (ADC) [44]. The symmetric distance computation (SDC) corresponds to the case when the query vector is also compressed:
$n=\underset{i=1..N}{\mathrm{argmin}}\|D(C(q))-D(C_{i})\|.$ (11)
@@ 4.1 The vector codecs
The ideal vector quantizer minimizes the MSE between the original and the decompressed vectors. This is formalized in the Lloyd necessary conditions for the optimality of a quantizer [52].
@@
A very classical and simple scalar quantizer is LSH (IndexLSH), where each vector component is encoded in a single bit by comparing it to a threshold. The threshold can be fixed to 0 or trained. Faiss supports efficient SDC search of binary vectors via the IndexBinary objects, see Section 4.5.
The ScalarQuantizer also supports uniform quantizers that encode a vector component into 8, 6 or 4 bits – referred to as SQ8, SQ6, SQ4. A scale and offset determine which values are reconstructed. They can be set separately for each dimension on the whole vector. The IndexRowwiseMinMax stores vectors with per-vector normalizing coefficients. Lower-precision 16-bit floating point representations are also considered as scalar quantizers, SQfp16 and SQbf16.
@@
The product quantizer (ProductQuantizer, also noted PQ) is a simple MCQ that splits the input vector into $M$ sub-vectors and quantizes them separately [44] with a k-means quantizer. At reconstruction time, the individual reconstructions are concatenated to produce the final code. In the following, we will use the notation PQ6x10 for a product quantizer with 6 sub-vectors each encoded in 10 bits ($M=6$, $K=2^{10}$).
Additive quantizers are a family of MCQ where the reconstructions from sub-quantizers are summed up together. Finding the optimal encoding for a vector given the codebooks is NP-hard [4], so, in practice, additive quantizers use heuristics to find near-optimal codes.
Faiss supports two types of additive quantizers. The residual quantizer (ResidualQuantizer) proceeds sequentially, by encoding the difference (residual) of the vector to encode and the one that is reconstructed by the previous sub-quantizers [19]. The local search quantizer (LocalSearchQuantizer) starts from a sub-optimal encoding of the vector and locally explores neighbording codes in a simulated annealing process [56, 57]. We use notations LSQ6x10 and RQ6x10 to refer to additive quantizers with 6 codebooks of size $2^{10}$.
Faiss also supports a combination of PQ and additive quantizer, ProductResidualQuantizer. In that case, the vector is split in sub-vectors that are encoded independently with additive quantizers [5]. The codes from the sub-quantizers are concatenated. We use the notation PRQ2x6x10 to indicate that vectors are split in 2 and encoded independently with RQ6x10, yielding a total of 12 codebooks of size $2^{10}$.
@@
Although this is not by design, there is a strict ordering between the quantizers described before. This means that quantizer $i+1$ can have the same set of reproduction values as quantizer $i$: it is more flexible and more data adaptive. The hierarchy of quantizers is:
@@
    the scalar quantizer is a product quantizer with 1 dimension per sub-vector and uniform per-dimension quantizer;
@@
    the product quantizer is a product-additive quantizer where the additive quantizer has a single level;
@@
    the product additive quantizer is an additive quantizer where within each codebook all components outside one sub-vector are set to 0 [4];
@@
    the additive quantizer is the general case where the codebook entries correspond to all possible reconstructions obtained by adding elements from the subquantizers.
The implications of this hierarchy are (1) the degrees of freedom for the reproduction values of quantizer $i+1$ are larger than for $i$, so it is more accurate (2) quantizer $i+1$ has a higher capacity so it consumes more resources in terms of training time and storage overhead than $i$. In practice, the product quantizer often offers a good trade-off, which explains its wide adoption. The corresponding Faiss Quantizer objects are listed in Appendix A.6.
Figure 2: Comparison of additive quantizers in terms of encoding time vs. accuracy (MSE). Lower values are better for both. We consider two different regimes: Deep1M (low-dimensional) to 8-bytes codes and Contriever1M (high dimensional) to 64-byte codes. For some RQ variants, we indicate the beam size setting at which that trade-off was obtained.
@@ 4.2 Vector preprocessing
Applying transformations to input vectors before encoding can enhance the effectiveness of certain quantizers. In particular, $d$-dimensional rotations are commonly used, as they preserve comparison metrics like cosine, L2 and inner product.
Scalar quantizers assign the same number of bits per vector component. However, for distance comparisons, if specific vector components have a higher variance, they have more impact on the distances. In other works, a variable number of bits are assigned per component [71]. However, it is simpler to apply a random rotation to the input vectors, which in Faiss can be done with a RandomRotationMatrix. The random rotation spreads the variance over all the dimensions without changing the measured distances.
An important transform is the Principal Component Analysis (PCA), that reduces the number of dimensions $d$ of the input vectors to a user-specified $d^{\prime}$. This operation (PCAMatrix) is the orthogonal linear mapping that best preserves the variance of the input distribution. It is often beneficial to apply PCA to large input vectors before quantizing them as k-means quantizers are more likely to “fall” in local minima in high-dimensional spaces [51, 45].
The OPQ transformation [33] is a rotation of the input space that decorrelates the distribution of each sub-vector of a product quantizer [^4] In Faiss terms, OPQ and ITQ are preprocessings. The actual quantization is performed by a subsequent product quantizer or binarization step.. This makes PQ more accurate in the case where the variance of the data is concentrated on a few components. The Faiss implementation OPQMatrix combines OPQ with a dimensionality reduction. The ITQ transformation [35] similarly rotates the input space prior to binarization (ITQMatrix).
@@ 4.3 Faiss additive quantization options
Additive quantizers exist in two main variants: the residual quantizer and local search quantizer. They are more complex than most quantizers because the index building time must be taken into account. In fact, the accuracy of an additive quantizer of a certain size can always be increased at the cost of an increased encoding time (and training time).
Additive quantizers are based on $M$ codebooks $T_{1},...T_{M}$ of size $K$. The decoding of code $C(x)=(c_{1},...,c_{M})$ is
$x^{\prime}=D(C(x))=T_{1}[c_{1}]+...+T_{M}[c_{M}].$ (12)
Thus, decoding is unambiguous. However, there is no practical way to encode vectors optimally, let alone train the codebooks. Enumerating all possible encodings is of exponential complexity in $M$.
@@
RQ encodes a vector $x$ sequentially. At stage $m$, RQ picks the entry that best reconstructs the residual of $x$ w.r.t. the previous encoding steps:
$c_{m}=\underset{j=1..K}{\mathrm{argmin}}\left\|\sum_{i=1}^{m-1}T_{i}[c_{i}]+T_{m}[j]-x\right\|^{2}.$ (13)
This greedy approach tends to get trapped in local minima. As a mitigation, the encoder maintains a beam of max_beam_size of possible codes and picks the best code at stage $M$. This parameter adjusts the trade-off between encoding time and accuracy.
@@
At encoding time, LSQ starts from a suboptimal encoding of the vector and proceeds with a simulated annealing optimization to refine the codes. At each optimization step, LSQ randomly flips codes and then uses Iterated Conditional Mode (ICM) to optimize the new encoding. The number of optimization steps is set with encode_ils_iters. The LSQ codebooks are trained via an expectation-maximization procedure (similar to k-means).
Deep1M Contriever1M
Refer to caption Refer to caption
Figure 3: Accuracy vs. code size trade-off for different codecs on the Deep1M and Contriever1M datasets. We show Pareto-optimal variants with larger dots and indicate the quantizer in text for some of them. Note that contriever vectors can be encoded to MSE=$2\cdot 10^{-4}$ in 768 bytes with SQ8 (that setting is widely out-of-range for the plot).
@@
$\langle q,x^{\prime}\rangle=\sum_{m=1}^{M}\langle T_{m}[c_{m}],q\rangle=\sum_{m=1}^{M}\mathrm{LUT}_{m}[c_{m}].$ (14)
The lookup tables $\mathrm{LUT}_{m}$ are computed when a query vector comes in, similar to product quantizer search [44].
This decomposition does not work to compute L2 distances. As a workaround, Faiss uses the decomposition [4]
$\|q-x^{\prime}\|^{2}=\|q\|^{2}+\|x^{\prime}\|^{2}-2\langle q,x^{\prime}\rangle.$ (15)
Thus, the term $\|x^{\prime}\|^{2}$ must be available at search time. Using the AdditiveQuantizer.search_type configuration, it can be appended in the stored code (ST_norm_float32), possibly compressed (ST_norm_qint8, ST_norm_qint4,…). It can also be computed on the fly (ST_norm_from_LUT) with
$\|x^{\prime}\|^{2}=2\sum_{m=1}^{M}\sum_{\ell=1}^{m-1}\langle T_{m}[c_{m}],T_{\ell}[c_{\ell}]\rangle+\sum_{m=1}^{M}\|T_{m}[c_{m}]\|^{2}.$ (16)
There, the norms and dot products are stored in the same lookup tables as the one used for beam search. Therefore, it trades off search time for memory overhead to store codes.
Figure 2 shows the trade-off between encoding time and MSE. Given a code size, it is more accurate to use a smaller number of sub-quantizers $M$ and a higher $K$. GPU encoding for LSQ does not help systematically. The LUT-based encoding of RQ is interesing for RQ/PRQ quantization when the beam size is larger. In the 64-byte regime, we observe that LSQ is not competitive with RQ. PLSQ and PRQ progressively become more competitive for larger memory budgets. They are also faster, since they operate on smaller vectors.